# Role of intermediate resonances in attosecond photoelectron interferometry in neon

### v1

@@ IV.2 Delay-resolved data
Figure 5: Top: RABBIT traces for photoelectron energies in the range $0-4$ eV for (a) 800 nm and (b) 1006 nm. Bottom: Variation of the sideband signal as a function of the relative delay between the IR and XUV fields, obtained by integrating over an energy window of $\approx 0.3$ eV for the lowest energy AT-RABBIT sideband (red circles) and the UT-RABBIT sideband (black squares) peaks at (c) 800 nm and (d) 1006 nm. The dotted lines are sinusoidal fits to the experimental data. The oscillations of the sideband signals are plotted relative to the average intensity of the sideband, which we set as the zero line for the normalized intensity.
We now focus the discussion on the comparison of the experimental and calculated data for the first sideband above the ionization threshold (UT-RABBIT). RABBIT traces as a function of the delay $\tau$ between the XUV and IR fields in the low kinetic energy region are shown for $\lambda=800$ nm and $\lambda=1006$ nm in Fig. 5(a) and (b), respectively. The intensity oscillations of the first two sidebands can be fitted by cosine functions (see Fig. 5(c,d)), from which the RABBIT phases can be extracted according to Eq. (3). The phase difference between the UT-RABBIT sideband and the first AT-RABBIT sideband is remarkably larger for the 800 nm wavelength driver ($\Delta\phi=2.39\pm 0.13$ rad) compared to the 1006 nm case ($\Delta\phi=-0.64\pm 0.27$ rad).
@@
Regarding $S_{2}$, there is better overall agreement among the three sets of theoretical predictions than for $S_{0}$, with the largest differences occurring in the near-threshold region (panel (c)). In the $1-7$ eV range (panel (d)), the theories agree well with each other and, after accounting for the different widths of the peaks, the agreement with the experimental predictions also improved compared to $S_{0}$.
Finally, the experimental RABBIT phase is well reproduced by all theories in the energy range 3-7 eV (panel (f)), for the peaks around 3.4 eV, 4.9 eV, and 6.4 eV.
@@
In the near-threshold region, all theories predict a doublet structure in $S_{2}$ (panel (c)), similar to that seen in $S_{0}$ in Fig. 4(a) and Fig. 7(a). Here the RE model is closest to the experimental data regarding the energy dependence. It should be pointed out that the data presented in Figs. 4 and 7 were taken under slightly different experimental conditions than those presented in Fig. 2. For this reason, care should be taken in evaluating the quality of the comparison. Note, in particular, that the height of the experimental sidebands in panel (b) is relatively low, suggesting that the IR intensity in the calculations was significantly higher than in the experiment. This issue is also clearly seen in panel (d), where the agreement between the theories is at a level of about a factor of 2, while the calculated oscillation amplitudes are much larger than in the experimental data.
@@ IV.3 Triple-differential data
In this section we present the energy- and angle-resolved RABBIT phase that we can extract from the experimental and simulated sideband oscillations according to Eq. (II). As indicated in the discussion of the angle-integrated phases in Figs. 6 and 7, the phase is rather insensitive to details in the peak heights, i.e., the particular IR intensity, as long as one remains in the (second-order) perturbative regime. The evolution of the RABBIT phase $\phi(\theta)$ over the entire range of angles for the energy interval $0-3.5$ eV is presented in Fig. 8 for the driving field with $\lambda=1006$ nm. Overall, there is very satisfactory agreement between the experimental map and that predicted by the RMT model, especially in light of the fact that SB18 in the RMT calculation is narrower than in the experiment. The key characteristics of the experimental data are also reproduced by the PT approach. The same is true for the RE map (not shown), which is very similar to that obtained with PT.
@@ V Conclusions
We have presented a joint experimental and theoretical investigation of attosecond photoelectron interferometry around the ionization threshold of neon. Experimental data were acquired using attosecond pulse trains driven by two different driving fields with wavelengths of $\lambda\approx 800$ nm and $\lambda\approx 1006$ nm. For the latter configuration, fine-tuning of the driving laser wavelength in the range $\rm 1001~{}nm\leq\lambda\leq 1019~{}nm$ made it possible to change the relative weight of closely-spaced intermediate resonances. The angular-resolved data indicate how the presence of resonances is affecting the photoelectron angular distributions of the sidebands close to the ionization threshold.
@@ Acknowledgments
The experimental part of this work was supported by the European Union’s Horizon 2020 project 641789 MEDEA. D.B. acknowledges support from the Swedish Research Council grant 2020-06384. G.S. acknowledges support by the Deutsche Forschungsgemeinschaft through Priority Program 1840 (QUTIF) and grant 429805582 (Project SA 3470/4-1). I.M. and G.S. acknowledge support by the BMBF project 05K19VF1, the Deutsche Forschungsgemeinschaft through the Research Training Group DynCAM (RTG 2717), and the Georg H. Endress Foundation. D.E. and G.S. also acknowledge funding by the DFG project International Research Training Group (IRTG) CoCo 2079 and INST 39/1079. The theoretical part of this work was supported by the NSF under OAC-1834740, PHY-2110023, PHY-2408484 (K.B.), OAC-2311928 (K.R.H. and K.B.), the ACCESS allocation PHY-090031, and the Frontera Pathways allocation PHY20028. The RMT calculations were performed on Stampede-2 and Frontera at the Texas Advanced Computing Center.

### v2

@@ IV.2 Delay-resolved data
We now focus the discussion on the comparison of the experimental and calculated data for the first sideband above the ionization threshold (UT-RABBIT). RABBIT traces as a function of the delay $\tau$ between the XUV and IR fields in the low kinetic energy region are shown for $\lambda=800$ nm and $\lambda=1006$ nm in Fig. 5(a) and (b), respectively. The intensity oscillations of the first two sidebands can be fitted by cosine functions (see Fig. 5(c,d)), from which the RABBIT phases can be extracted according to Eq. (3). The phase difference between the UT-RABBIT sideband and the first AT-RABBIT sideband is remarkably larger for the 800 nm wavelength driver ($\Delta\phi=2.39\pm 0.16$ rad) compared to the 1006 nm case ($\Delta\phi=-0.65\pm 0.27$ rad). The large phase shift for the first wavelength is well reproduced by all the simulation models, as shown in Fig. 5, although with some difference in the exact value (see Table 2). The phase shift for the $1006$-nm-case is partially reproduced only by the RMT simulations, while the PT and RE model predict a small difference but with the opposite sign to the experimental one. We attribute the differences observed for the three different models in the second case to the description of the large number of states below the ionization threshold.
Figure 5: RABBIT traces (a,b,e,f,i,j,m,n) in the range $0-4$ eV and oscillations of the UT-RABBIT (black line) and first AT-RABBIT (red lines) sidebands (panels c,d,g,h,k,l,o,p) extracted from the experimental data (a-d), RMT (e-h), PT (i-l), and RE (m-p) models for $\lambda=800$ nm (left panels) and $\lambda=$1006 nm (right panels). The oscillations of the sideband signals are plotted relative to the average intensity of the sideband, which we set as the zero line for the normalized intensity. The dotted lines in panels (a) and (b) are sinusoidal fits to the experimental data.
$\lambda=800$ nm
$\phi(\mathrm{UT})$ (rad) $\phi(\mathrm{AT})$ (rad) $\Delta\phi$ (rad)
Exp. $3.46\pm 0.08$ $1.07\pm 0.08$ $2.39\pm 0.16$
RMT $4.25$ $1.04$ $3.21$
PT $3.72$ $1.19$ $2.53$
RE $4.21$ $1.32$ $2.89$
$\lambda=1006$ nm
$\phi(\mathrm{UT})$ (rad) $\phi(\mathrm{AT})$ (rad) $\Delta\phi$ (rad)
Exp. $0.12\pm 0.15$ $0.77\pm 0.12$ $-0.65\pm 0.27$
RMT $-0.92$ $0.39$ $-1.31$
PT $0.97$ $0.50$ $0.47$
RE $1.07$ $0.53$ $0.54$
Table 2: Phase $\phi$ extracted using sinusoidal fits from the experimental data and simulations using the three different models for the UT and AT sidebands and for the two driving wavelengths $\lambda=800$ nm and $\lambda=1006$ nm. The phase difference $\Delta\phi=\phi(\mathrm{UT})-\phi(\mathrm{AT})$ is also reported. The error bars for the experimental data derived directly from the fitting algorithm.
@@
Regarding $S_{2}$, there is better overall agreement among the three sets of theoretical predictions than for $S_{0}$, with the largest differences occurring in the near-threshold region (panel (c)). In the $1-7$ eV range (panel (d)), the theories agree well with each other and, after accounting for the different widths of the peaks, the agreement with the experimental predictions also improved compared to $S_{0}$. Finally, the experimental RABBIT phase is well reproduced by all theories in the energy range 3-7 eV (panel (f)), for the peaks around 3.4 eV, 4.9 eV, and 6.4 eV.
@@
In the near-threshold region, all theories predict a doublet structure in $S_{2}$ (panel (c)), similar to that seen in $S_{0}$ in Fig. 4(a) and Fig. 7(a). Here the RE model is closest to the experimental data regarding the energy dependence. We attribute this agreement to the description of the fine-structure splitting. Even though RMT is an ab-initio method, it describes the process in the LS-scheme. In our case, however, the 3d states are highly mixed due to the fine-structure interaction. This is best described in the jK-coupling scheme adopted in the RE approach. It should be pointed out that the data presented in Figs. 4 and 7 were taken under slightly different experimental conditions than those presented in Fig. 2. For this reason, care should be taken in evaluating the quality of the comparison. Note, in particular, that the height of the experimental sidebands in panel (b) is relatively low, suggesting that the IR intensity in the calculations was significantly higher than in the experiment. This issue is also clearly seen in panel (d), where the agreement between the theories is at a level of about a factor of 2, while the calculated oscillation amplitudes are much larger than in the experimental data.
@@ IV.3 Triple-differential data
In this section we present the energy- and angle-resolved RABBIT phase that we can extract from the experimental and simulated sideband oscillations according to Eq. (1). As indicated in the discussion of the angle-integrated phases in Figs. 6 and 7, the phase is rather insensitive to details in the peak heights, i.e., the particular IR intensity, as long as one remains in the (second-order) perturbative regime. The evolution of the RABBIT phase $\phi(\theta)$ over the entire range of angles for the energy interval $0-3.5$ eV is presented in Fig. 8 for the driving field with $\lambda=1006$ nm. Overall, there is very satisfactory agreement between the experimental map and that predicted by the RMT model, especially in light of the fact that SB18 in the RMT calculation is narrower than in the experiment. The key characteristics of the experimental data are also reproduced by the PT approach. The same is true for the RE map (not shown), which is very similar to that obtained with PT.
@@ V Conclusions
We have presented a joint experimental and theoretical investigation of attosecond photoelectron interferometry around the ionization threshold of neon. Experimental data were acquired using attosecond pulse trains driven by two different driving fields with wavelengths of $\lambda\approx 800$ nm and $\lambda\approx 1006$ nm. For the latter configuration, fine-tuning of the driving laser wavelength in the range $\rm 1001~{}nm\leq\lambda\leq 1019~{}nm$ made it possible to change the relative weight of closely-spaced intermediate resonances. The angular-resolved data indicate how the presence of resonances is affecting the photoelectron angular distributions of the sidebands close to the ionization threshold. The presence of resonances below the ionization threshold can affect the photoionization phase in RABBIT measurements depending on the specific resonance and photon energy. In the case of atoms, the contributions of the different resonances can be disentangled by exploiting the tunability of the laser and the resolution of the photoelectron spectrometer. In molecular systems, isolating the contributions of the different resonances is a major challenge due to the complex electronic and vibrational structure. Their influence on the RABBIT spectra will require the development of accurate theoretical methods [45, 46] to describe the dense energy distribution around the ionization threshold.
@@ Acknowledgments
The experimental part of this work was supported by the European Union’s Horizon 2020 project 641789 MEDEA. D.B. acknowledges support from the Swedish Research Council grant 2020-06384. G.S. acknowledges support by FRIAS. I.M. and G.S. acknowledge support by the BMBF project 05K19VF1 and the Georg H. Endress Foundation. The theoretical part of this work was supported by the NSF under OAC-1834740, PHY-2110023, PHY-2408484 (K.B.), OAC-2311928 (K.R.H. and K.B.), the ACCESS allocation PHY-090031, and the Frontera Pathways allocation PHY20028. The RMT calculations were performed on Stampede-2 and Frontera at the Texas Advanced Computing Center.